2. Citizens of the country of Heehaw produce hay and provide entertainment services (banjo playing). In one year they produced $15 million worth of hay, with S11 million consumed domestically and the other $4 million sold to neighboring countries. They provided $7 million worth of banjo-playing services, $5 million in Heehaw, and $2 million in neighboring countries. They purchased $6 million worth of soda pop from neighboring countries. Calculate the magnitudes of GNP, GDP, net factor payments from abroad, net exports, and the current account balance.
Answer #1

GNP = Total goods produced in a nation = 15+7 = $22 million
GDP = goods produced in domestic = hay produced domestically+ domestic banjo playing = 15+5=$20 million
Net factor payments = GNP-GDP = 22-20 = $2 million
Net export = hay sold abroad-soda pop imports = 4-6 = -$2 million
Current account balance = net exports + net factor payments = -2+2 = $0 million